# Transforming Strategic Games into Biform Games: Applications in Allocation Mechanisms and Green Technology Investment

### v1

@@ 2.3 Distribution of Cooperative Outcomes and Biform Game Solutions
Corollary 2.1 For a given $\left({N,{u_{f}},X}\right)$, if ${a^{{u_{f}}}}\left(x\right)$ is the Shapley allocation function of $\left({N,{u_{f}}\left({\cdot,x}\right)}\right)$, then $x$ is a solution of $\left({N,{v_{f}},X}\right)$ if and only if $x$ is a Nash equilibrium of $\left({N,\left({{X_{i}}}\right),\left({a_{i}^{{v_{f}}}}\right)}\right)$.
Proposition 2.1 and Corollary 2.1 show that a marginalist distribution method cannot promote mutually beneficial cooperation. In this case, the solution to the biform game derived from the strategic game theory is still the non-cooperative solution to the game, namely, the Nash equilibrium of $\left({N,\left({{X_{i}}}\right),\left({{f_{i}}}\right)}\right)$. This shows that if the cooperative outcomes are distributed strictly according to marginalism, the biform game derived from the strategic game theory will still be in a non-cooperative state.
Proposition 2.2 If the allocation function $a^{{v_{f}}}$} is egalitarian, then the strategy combination ${x^{*}}\in\mathop{argmax}\limits_{x\in X}{v_{f}}\left({N,x}\right)$ is a Nash equilibrium of
$\left({N,\left({{X_{i}}}\right),\left({{f_{i}}}\right)}\right)$, and thus is also a solution of $\left({N,{v_{f}},X}\right)$.
@@
According to Proposition 2.2, the egalitarian distribution method will maximize the grand coalition’s benefit ${v_{f}}\left({N,x}\right)$, that is, the maximum point of ${v_{f}}\left({N,x}\right)$ becomes the solution to the biform game, which shows that the egalitarian distribution method will promote cooperation and achieve the maximum overall benefit.
Next, we further discuss the case where there is no additional benefit, that is, the case where ${v_{f}}={u_{f}}$.
@@
then $f\left(x\right)=\left({{f_{1}}\left(x\right),\cdots,{f_{n}}\left(x\right)}\right)$ is said to be the Pareto optimal payoff of $\left({N,X,f}\right)$, or the strategy combination $x\in X$ is said to be Pareto optimal. In other words, if a Pareto optimal outcome is to be improved, the interests of at least one participant will inevitably be harmed. Usually, the Nash equilibrium is not Pareto optimal, which means that the benefits of all players can be improved.
@@
Corollary 2.2 If the allocation function $a^{{v_{f}}}$} is egalitarian, then the strategy combination ${x^{*}}\in\mathop{argmax}\limits_{x\in X}{v_{f}}\left({N,x}\right)$ is a solution to $\left({N,{u_{f}},X}\right)$, and $f\left({{x^{*}}}\right)=\left({{f_{1}}\left({{x^{*}}}\right),\cdots,{f_{n}}\left({{x^{*}}}\right)}\right)$ is the Pareto optimal payoff of $\left({N,X,f}\right)$.
The above results show that egalitarian distribution promotes cooperation and is conducive to achieving Pareto optimality. Marginalism, while emphasizing the contributions of players, is not conducive to cooperation, as the ultimate outcome still reverts to non-cooperation. These conclusions not only reveal the nature of the solutions to biform games but also offer a new perspective on the conceptual study of cooperative game solutions. It can also be argued that they offer a new perspective on the distribution of any type of cooperation and shared gains.
@@ 2.4 Give examples to illustrate
Example 2.2 On a common pasture, two herders raise sheep. Let ${q_{1}},{q_{2}}$ denote each herder’s sheep stock, and  ${\text{\;}}{\rho_{1}},{\text{\;}}{\rho_{2}}$ denote their profits from raising sheep. Given $M>0$, $M$ denotes the carrying capacity of the pasture for sheep. That is, if the total sheep stock is ${q^{+}}={q_{1}}+{q_{2}}>M$, then each herder’s profit  ${\text{\;}}{\rho_{i}}<0$. Therefore, we assume that the herders’ sheep stock ${q_{i}}\in\left[{0,M}\right]$. Let $\mu$ denote the slaughter rate per sheep on the pasture, which is the ratio of the total sheep stock $q^{+}$}. Assume that $\mu\left({{q^{+}}}\right)$ is a second-order differentiable, decreasing, concave function with respect to $q^{+}$}. Let $p_{0}$} denote the unit price of mutton, and $c_{0}$} denote the cost of raising a sheep for each herder (including labor input).
@@
Let ${X_{1}}={X_{2}}=\left[{0,M}\right]$, $q=\left({{q_{1}},{q_{2}}}\right)$, and $\rho=\left({{\rho_{1}},{\rho_{2}}}\right)$, then $\left({\left\{{1,2}\right\},X,\rho}\right)$ constitute a strategic game, where $\rho_{1}$} and $\rho_{2}$} are the payoff functions of herders 1 and 2. For ease of discussion, we assume that the number of sheep on hand is considered a continuous variable, the unit price $p_{0}$} is set to 1, and $0\leqslant{c_{0}}\leqslant 1$ is considered a constant.
@@
According to the assumption of $M$, when ${q_{1}}=M$, ${\rho_{1}}\leqslant 0$; when ${q_{2}}=M$, ${\rho_{2}}\leqslant 0$. Therefore, for every ${q_{2}}\in\left[{0,M}\right]$, the maximum value of $\rho_{1}$} can only be achieved within $\left({0,M}\right)$; similarly, for every ${q_{1}}\in\left[{0,M}\right]$, the maximum value of $\rho_{2}$} can only be achieved within $\left({0,M}\right)$.
@@
It is easy to see that ${q_{1}}={q_{2}}$. So the above equations are equivalent to the following equation
@@
By the decreasing and concave nature of $\mu\left({{q^{+}}}\right)$, we know that the above equation has a unique solution, denoted as $q^{+*}$}. Then, by ${q_{1}}={q_{2}}=\frac{{{q^{+*}}}}{2}$, we can obtain the Nash equilibrium of $\left({\left\{{1,2}\right\},\left({{X_{i}}}\right),\left({{\rho_{i}}}\right)}\right)$ as $\left({\frac{{{q^{+*}}}}{2},\frac{{{q^{+*}}}}{2}}\right)$. At this point, the profits of herders 1 and 2 are both
@@
Similar to the previous method, from the decreasing and concave properties of $\mu\left({{q^{+}}}\right)$, it can be seen that the only maximum point of $\rho$ is the solution of the following equation:
@@
Since $\mu\left({{q^{+}}}\right)$ is a decreasing concave function and $q^{+\#}$} is the maximum point of $\rho$, then $\rho^{\prime}\left({{q^{+\#}}}\right)=0$, and when ${q^{+}}\geqslant{q^{+\#}}$, $\rho^{\prime}\left({{q^{+}}}\right)\geqslant 0$. Therefore, ${q^{+\#}}<{q^{+*}}$.
Next, we compare the profits of the two cases. Notice
$\begin{split}\rho(q^{+\#})-\rho(q^{+*})&=\left[\mu(q^{+\#})q^{+\#}-q^{+\#}c_{0}\right]-\left[\rho_{1}\left(\frac{q^{*}}{2}\right)+\rho_{2}\left(\frac{q^{*}}{2}\right)\right]\\ &>\left[\mu(q^{+\#})q^{+\#}-q^{+\#}c_{0}\right]-\left[\mu(q^{+*})q^{+*}-q^{+*}c_{0}\right]\\ &>\left[\mu(q^{+\#})q^{+\#}-\mu(q^{+*})q^{+*}\right]+\left[q^{+*}c_{0}-q^{+\#}c_{0}\right].\end{split}$
Therefore, there exists $\zeta\in$ $\left({{q^{+\#}},{q^{+*}}}\right)$ such that
$\rho(q^{+\#})-\rho(q^{+*})>\mu^{\prime}(\zeta)(q^{+\#}-q^{+*})+c_{0}(q^{+*}-q^{+\#})=(c_{0}-\mu^{\prime}(\zeta))(q^{+*}-q^{+\#})>0.$
Therefore, $\rho\left({{q^{+\#}}}\right)>\rho\left({{q^{+*}}}\right)$. This shows that the total profit corresponding to the Nash equilibrium $(\frac{{{q^{+*}}}}{2},\frac{{{q^{+*}}}}{2}$) is lower than the case of raising fewer sheep, that is, the total profit is lower than the case of $\left({\frac{{{q^{+\#}}}}{2},\frac{{{q^{+\#}}}}{2}}\right)$.
We consider this problem as a biform game, and the coalition function ${v_{\rho}}:{2^{N}}\times X\to R$ associated with the strategy combination $x$. Here we only need to consider the case $v_{\rho}(S,\cdot)=u_{\rho}(S,\cdot)=\sum_{i\in S}\rho_{i}(\cdot)$ where there is no additional benefit. The coalition function is as follows:
$\begin{split}v_{\rho}(\emptyset,(q_{1},q_{2}))&=0,\\ v_{\rho}(\{1\},(q_{1},q_{2}))&=u_{\rho}(\{1\},(q_{1},q_{2}))=\rho_{1}=\mu(q)q_{1}-q_{1}c_{0},\\ v_{\rho}(\{2\},(q_{1},q_{2}))&=u_{\rho}(\{2\},(q_{1},q_{2}))=\rho_{2}=\mu(q)q_{2}-q_{2}c_{0},\\ v_{\rho}(\{1,2\},(q_{1},q_{2}))&=u_{\rho}(\{1,2\},(q_{1},q_{2}))=\rho_{1}+\rho_{2}=\mu(q)-qc_{0}.\end{split}$

### v2

@@ 2.3 Distribution of Cooperative Outcomes and Biform Game Solutions
Corollary 2.1 For a given $\left({N,{u_{f}},X}\right)$, if ${a^{{u_{f}}}}\left(x\right)$ is the Shapley allocation function of $\left({N,{u_{f}}\left({\cdot,x}\right)}\right)$, then $x$ is a solution of $\left({N,{u_{f}},X}\right)$ if and only if $x$ is a Nash equilibrium of $\left({N,\left({{X_{i}}}\right),\left({u_{f}}\right)}\right)$.
Proposition 2.1 and Corollary 2.1 show that a marginalist distribution method cannot promote mutually beneficial cooperation. In this case, the solution to the biform game derived from the strategic game theory is still the non-cooperative solution to the game, namely, the Nash equilibrium of $\left({N,\left({{X_{i}}}\right),\left({{f_{i}}}\right)}\right)$. This shows that if the cooperative outcomes are distributed strictly according to marginalist, the biform game derived from the strategic game will still be in a non-cooperative state.
Proposition 2.2 If the allocation function $a^{{v_{f}}}$} is egalitarian, then the strategy profile ${x^{*}}\in\mathop{argmax}\limits_{x\in X}{v_{f}}\left({N,x}\right)$ is a Nash equilibrium of $\left({N,\left({{X_{i}}}\right),\left({{f_{i}}}\right)}\right)$, and thus is also a solution of $\left({N,{v_{f}},X}\right)$.
@@
According to Proposition 2.2, the egalitarian distribution will maximize the grand coalition’s benefit ${v_{f}}\left({N,x}\right)$, that is, the maximum point of ${v_{f}}\left({N,x}\right)$ becomes the solution to the biform game, which shows that the egalitarian distribution will promote cooperation and achieve the maximum overall benefit.
Next, we further discuss the case where there is no extra benefit, that is, the case where ${v_{f}}={u_{f}}$.
@@
then $f\left(x\right)=\left({{f_{1}}\left(x\right),\cdots,{f_{n}}\left(x\right)}\right)$ is said to be the Pareto optimal payoff of $\left({N,X,f}\right)$, or the strategy profile $x\in X$ is said to be Pareto optimal. In other words, if a Pareto optimal outcome is to be improved, the interests of at least one participant will inevitably be harmed. Usually, the Nash equilibrium is not Pareto optimal, which means that the benefits of some players can be improved.
@@
Corollary 2.2 If the allocation function $u_{f}$ is egalitarian, then the strategy profile ${x^{*}}\in\mathop{argmax}\limits_{x\in X}{u_{f}}\left({N,x}\right)$ is a solution to $\left({N,{u_{f}},X}\right)$, and $f\left({{x^{*}}}\right)=\left({{f_{1}}\left({{x^{*}}}\right),\cdots,{f_{n}}\left({{x^{*}}}\right)}\right)$ is the Pareto optimal payoff of $\left({N,X,f}\right)$.
The above results show that egalitarian distribution increas the willingness to cooperate and is conducive to achieving Pareto optimality. Marginalism, while emphasizing the contributions of players, is not conducive to cooperation, as the ultimate outcome still reverts to non-cooperation. These conclusions not only reveal the nature of the solutions to biform games but also offer a new perspective on the conceptual study of cooperative game solutions. It can also be argued that they offer a new perspective on the distribution of any type of cooperation and shared gains.
@@ 2.4 Give examples to illustrate
Example 2.2 On a common pasture, two herders raise sheep. Let ${q_{1}},{q_{2}}$ denote each herder’s sheep stock, and  ${\text{\;}}{\rho_{1}},{\text{\;}}{\rho_{2}}$ denote their profits from raising sheep. Given $M>0$, $M$ denotes the carrying capacity of the pasture for sheep. That is, if the total sheep stock is ${q^{+}}={q_{1}}+{q_{2}}>M$, then each herder’s profit  ${\text{\;}}{\rho_{i}}<0$. Therefore, we assume that the herders’ sheep stock ${q_{i}}\in\left[{0,M}\right]$. Let $\mu$ denote the slaughter rate per sheep on the pasture, which is the ratio of the total sheep stock $q^{+}$}. Assume that $\mu(M){p_{0}}={c_{0}}$, which means that if the total number of sheep raised exceeds the carrying capacity, there will be no profit at all, and $\mu\left({{q^{+}}}\right)$ is a second-order differentiable, decreasing, concave function with respect to $q^{+}$}. Let $p_{0}$} denote the unit price of mutton, and $c_{0}$} denote the cost of raising a sheep for each herder (including labor input).
@@
Let $X={X_{1}}\times{X_{2}}=\left[{0,M}\right]\times\left[{0,M}\right]$, $q=\left({{q_{1}},{q_{2}}}\right)$, and $\rho=\left({{\rho_{1}},{\rho_{2}}}\right)$, then $\left({\left\{{1,2}\right\},X,\rho}\right)$ constitute a strategic game, where $\rho_{1}$} and $\rho_{2}$} are the payoff functions of herders 1 and 2. For ease of discussion, we assume that the number of sheep on hand is considered a continuous variable, the unit price $p_{0}$} is set to 1, and $0\leqslant{c_{0}}\leqslant 1$ is considered a constant.Then the profit function is simplified as:
$\begin{split}\rho_{1}=\mu\left({{q^{+}}}\right){q_{1}}-{q_{1}}{c_{0}},\\ \rho_{2}=\mu\left({{q^{+}}}\right){q_{2}}-{q_{2}}{c_{0}}.\end{split}$
@@
It is not difficult to verify that the solution to the above equation satisfying ${q_{1}}={q_{2}}$. So the above equations are equivalent to the following equation
@@
Noting the decreasing and concave nature of $\mu\left({{q^{+}}}\right)$, $\mu(0)>c_{0}$ and $\mu(M)=c_{0}$, we know that the above equation has a unique solution, denoted as $q^{+*}$} which lies in $(0,M)$. Then, by ${q_{1}}={q_{2}}=\frac{{{q^{+*}}}}{2}$, we can obtain the Nash equilibrium of $\left({\left\{{1,2}\right\},\left({{X_{i}}}\right),\left({{\rho_{i}}}\right)}\right)$ as $\left({\frac{{{q^{+*}}}}{2},\frac{{{q^{+*}}}}{2}}\right)$. At this point, the profits of herders 1 and 2 are both
@@
Similar to the previous method, from the decreasing and concave properties of $\mu\left({{q^{+}}}\right)$, it can be seen that the unique maximum point of $\rho$ is the solution of the following equation:
@@
Since $\mu\left({{q^{+}}}\right)$ is a decreasing concave function and $q^{+\#}$} is the maximum point of $\rho$, then $\rho^{\prime}\left({{q^{+\#}}}\right)=0$, and when ${q^{+}}\geqslant{q^{+\#}}$, then $\rho^{\prime}\left({{q^{+}}}\right)<0$. Therefore, ${q^{+\#}}<{q^{+*}}$, and $\rho\left({{q^{+\#}}}\right)>\rho\left({{q^{+*}}}\right)$.
This shows that the total profit corresponding to the Nash equilibrium $(\frac{{{q^{+*}}}}{2},\frac{{{q^{+*}}}}{2}$) is lower than the case of raising fewer sheep, that is, the total profit is lower than the case of $\left({\frac{{{q^{+\#}}}}{2},\frac{{{q^{+\#}}}}{2}}\right)$.
A biform game is from the strategy game and the coalition function ${v_{\rho}}:{2^{N}}\times X\to R$ associated with the strategy profile $x$. Here we only need to consider the case $v_{\rho}(S,\cdot)=u_{\rho}(S,\cdot)=\sum_{i\in S}\rho_{i}(\cdot)$ where there is no extra benefit. The coalition function is as follows:
$\begin{split}v_{\rho}(\emptyset,(q_{1},q_{2}))&=0,\\ v_{\rho}(\{1\},(q_{1},q_{2}))&=u_{\rho}(\{1\},(q_{1},q_{2}))=\rho_{1}=\mu(q^{+})q_{1}-q_{1}c_{0},\\ v_{\rho}(\{2\},(q_{1},q_{2}))&=u_{\rho}(\{2\},(q_{1},q_{2}))=\rho_{2}=\mu(q^{+})q_{2}-q_{2}c_{0},\\ v_{\rho}(\{1,2\},(q_{1},q_{2}))&=u_{\rho}(\{1,2\},(q_{1},q_{2}))=\rho_{1}+\rho_{2}=\mu(q^{+})-qc_{0}.\end{split}$